Mike, Mark, and Melvin painted a fence. Mike painted the first 1/3 of the fence. Mark painted 3/5 of what remained. Melvin painted the final 40ft. How long is the fence?
1
Expert's answer
2010-07-08T04:41:54-0400
Let the length of the fence is x, so Mike painted (x/3), Mark painted ((x-x/3)*3/5) and Melvin painted 40. We can write the equation:

(х/3)+((х-(х/3))*3/5)+40 = х
(х/3)+(2х/5)+40 = х
х-(х/3)-(2х/5) = 40
4х/15 = 40
х = 150

So, the length of the fence is 150 ft.